OpenStudy (anonymous):

•the cube root of x cubed •:x to the one–third power • :x to the one–third power • x to the one–third power •1 over x to the –1 power •the eleventh root of the quantity of x to the fifth times x to the fourth times x squared

OpenStudy (anonymous):

\[3\sqrt[]{x3}\] \[x1/3 \times x1/3\times x1/3\] \[1/X ^-3\] \[11\sqrt[]{x}5\]\[11\sqrt[?]{x}5 \times x4 \times x2\]

OpenStudy (anonymous):

I need help with the names of the properties.